# Edge Graph Intelligence: Reciprocally Empowering Edge Networks with Graph Intelligence

### v1

@@ VIII Open Challenges and Future Research Opportunities
VIII-A New Promising Applications and Optimizations
@@
VIII-A1 B5G and 6G Protocols Design
@@
Communications can frequently occur in edge networks for various cross-device services. Their networking topology, identified by the communication nodes, is exactly a graph and all related data, such as channel capacity and queuing delays, can be modeled as graph data. This pertains to the opportunity to apply GI for developing advanced communication protocols, e.g., B5G and 6G, to enable data transmission in a smarter, faster, and greener way. For instance, one may use GNNs to dynamically schedule routing strategies and detect intermittent anomalies within edge networks [359, 360].
VIII-A2 Integrated Intelligent Applications
With increasingly more sensors and computing facilities universally embedded in smart cities and factories, massive amounts of streaming data are generated continuously. Unlike traditional methods that process these data solely from individual sensors, EGI collectively assembles them in a graph and thus can intelligently analyze them holistically with undiscovered insights. This rationale has driven existing services such as optimizing traffic flow, predicting energy consumption patterns, and identifying areas prone to crime, but is yet calling for more applications with graph data input, e.g., efficient resource allocation for public services and city-wide autonomous driving for Internet of Vehicles.
VIII-A3 Personalized Services
@@
End users nowadays are immersively surrounded by a vast amount of IoT devices like mobile smartphones, wearable kits, and smart home assets. For individual users, we can weave a user-centric network by collecting all user-related data from these surrounding devices and connecting them in a logic graph, on which EGI techniques can be applied to process, understand, and learn. In other words, EGI can serve as a tool to summarize data points scattered around individual users and be leveraged as a core component to build everyone-centric customized services.
VIII-B Edge-Centric Graph Learning Models and Systems
@@
Many existing GI models are originally designed to improve accuracy on specific tasks. In real deployment, however, a set of SLOs beyond accuracy are required by service providers such as responsive latency, system energy, and data privacy. Nonetheless, designing GI models tailored to edge services is non-trivial due to the unique characteristics and constraints of GI models and edge networks.
VIII-B1 Resource Constraints
@@
Edge devices typically have limited but heterogeneous computational capabilities compared to cloud servers, possessing platforms from MCUs to mobile smartphones to edge servers. Edge-centric GI models need to be lightweight, have acceptable memory footprints, and minimize computational complexity while still delivering accurate results, especially for EGI-based services with real-time or near real-time requirements.
VIII-B2 Resource Dynamics
@@
Edge devices frequently operate in dynamic and unpredictable resource conditions, including intermittent connectivity, dynamic computing capabilities, or fluctuated communication bandwidths. A stable, satisfactory EGI system is obliged to be robust to these variations and capable of adapting to different network conditions to ensure reliable and timely analysis, especially for those services that process graph data across multiple edge devices.
VIII-B3 Scalability
@@
Many real-world networks represented as graphs can be massive and evolve in size, but how to scale EGI techniques to large, dynamic graphs poses significant challenges on both the algorithm and system sides. From the model perspective, large graph models borrow wisdom from large language models and incorporate pre-training methods to embrace graphs on giant scales, but also bring significant challenges to EGI system design edge networks. Distributed edge collaboration (cf. Sec. V-C), which aggregates multiple edge devices as a resource pool for model computation, can be a potential solution but still requires tailored designs with parallel processing, graph partitioning, and incremental learning, etc.
In summary, addressing these aspects requires the development of innovative algorithms and optimization strategies specifically tailored for EGI.
VIII-C Emerging Learning Paradigms
@@
VIII-C1 Large Graph Model
@@
Large Graph Models (LGMs), or Graph Foundation Models (GFMs), borrow the wisdom of Large Language Models (LLMs) to the graph domain that simultaneously scales both training data and models. This scaling mechanism envisions LGMs to be capable of capturing complex relationships and patterns within large-scale graphs, which are common representations of data in various domains such as social networks, biological networks, and transportation systems. Nonetheless, applying LGM in edge networks severely suffers from the constrained computing capability and network capacity. Edge-cloud collaboration may be a potential solution for deployment, which marries the local data processing ability of edge networks and the powerful computing power of the cloud. Given the promising ability of large models, LGMs desire more effort in the context of EGI.
VIII-C2 Few-Shot Learning
@@
Few-shot learning means training GI models with a few labeled examples per class or category in a graph dataset. Its goal is to enable the GI model to generalize and make accurate predictions on unseen classes or categories with limited labeled data. This is particularly useful when edge devices collect a volume of data but only label a portion of them. For example, in some EGI-based object detection tasks, GI models can learn to recognize novel objects with a few labeled examples by leveraging information from similar objects and their relationships in a graph representation.
VIII-C3 Transfer Learning
Transfer learning for GNNs leverages knowledge learned from a source domain or task to improve performance in a target domain or task. In other words, the pre-trained GI model is used as a starting point and fine-tuned on the target domain with limited labeled data. Transfer learning empowers GI models with the ability to adapt their learned distribution, which is relevant in many edge applications since data collected in edge environments are usually dynamically changing. An instance of transfer graph learning is anomaly detection in networked systems, where GI can learn patterns of fraud by pre-training on a large graph and transfer this knowledge to identify fraudulent activities in a target domain with limited labeled data.
VIII-C4 Continual Learning
@@
Continual Learning (CL) for GI is to incrementally retrain GNN models to adapt to new data or tasks without forgetting previous knowledge. It also adjusts the learned latent distribution like Transfer Learning but emphasizes learning patterns from the additional incoming data while preserving previously learned representations. For EGI services, CL enables incremental service refinement for evolved graphs, e.g., in temporal graph data analysis tasks, without training GI models from scratch. Examples include predicting traffic conditions in a road network and running status of power grids, where continual learning of GI models allow the model to adapt over time and incorporate new temporal information while retaining the knowledge learned from past time steps.
VIII-D Native Edge Support for Graph Intelligence
@@
VIII-D1 Graph Data Collection
@@
EGI applications can not be realized without pervasive graph data. For many applications, however, these graph data are generated and distributed in different places, e.g., traffic sensory data scattered in road networks and perception information gathered by vehicles in robotic swarms. How to collect these (geographically) distributed data over the edge networks efficiently, i.e., in a high-quality, low-latency, and privacy-preserved manner, thereby becomes an unavoidable problem in running GI-based services.
VIII-D2 Edge-Cloud Collaboration
@@
Edge networks render computing power in versatile forms, e.g., by a single edge device, a pool of collaborated ones, and even a synergetic edge-cloud continuum. Provisioning these versatile resources efficiently for EGI applications not only requires flexible resource management but also judicious workload scheduling that well aligns computation and communication. In particular, edge-cloud collaboration provides a combined mechanism to embrace both data locality at the edge and resource richness in the cloud, which allows GI to be more efficient and sustainable and attracts growing attention from the community. Possible means towards this objective include designing GI-oriented communication protocols, GI-specific computation scheduling mechanisms, and a holistic edge network cost optimization methodology.
VIII-D3 System Development and Deployment
@@
Given the broad use cases of EGI, their development yet lacks a comprehensive framework with full-stack toolkits, particularly for distributed edge collaboration and edge-cloud synergy. This requires addressing the complexity of distributed systems, including edge device coordination, data synchronization, and fault tolerance, and also tackling long-term interoperability and testing issues. Existing virtualization techniques such as container and function computing can be further explored for friendly EGI development and deployment.
Overall, concerning the lifecycle of EGI services, there are plentiful opportunities for building EGI-native edge infrastructures.
VIII-E Incentive Mechanisms and Business Models
The interaction between GI and edge networks fosters EGI as a digital flywheel, which drives a flourishing ecosystem that encompasses various entities such as GI model providers, edge device providers, network operators, data generators, and service consumers. The seamless operation of high-efficiency EGI services necessitates close collaboration and integration among different entities, e.g., for facilitating trustworthy resource sharing and smooth service handover. Thus, the establishment of an appropriate incentive mechanism and business model becomes indispensable to cultivating effective and efficient cooperation among the EGI ecosystem’s members. Specifically, from the perspective of edge networks, designing incentive mechanisms that encourage the participation of edge devices and fair resource sharing among them is crucial to initialize a business centering around EGI. Such encouragement can be achieved by proper reward schemes, and fairness should consider factors such as resource capacity, contribution level, and the value associated. From the perspective of GI models, a user may act as both an API consumer for model inference and a data generator for model training, and thus smart pricing schemes are imperative to appropriately assess the user’s service consumption and the value contributed by their data.
For EGI applications with decentralized graph data sources or decentralized computing facilities, blockchain and smart contract techniques may be incorporated into their procedures to examine the contribution, charge the price, and allocate revenue among pipeline members. A lightweight, EGI-oriented blockchain consensus protocol is also highly desirable for EGI consortiums.
VIII-F Explainability, Security, and Privacy
@@
The open nature of EGI requires trustworthiness across different entities in the ecosystem to make EGI happen. Nonetheless, in many scenarios, edge services are practical only they are reliable. For EGI, this relates to explainability, security, and privacy.
VIII-F1 Explainability
@@
VIII-F2 Security
Security has dual means for EGI services. On the one hand, GI models should be secure and robust since they are vulnerable if there are malicious vertices in the graph input. When the input graph data are collected from multiple entities, the risk of encountering fraud becomes remarkably high, which further highlights the necessity of building and applying secure GI models. On the other hand, edge networks also suffer from malicious devices that can disrupt the learning process or compromise the integrity of the data. This demands building verifiable trust mechanisms among edge networks and developing secure tools to defend and recover from these attacks.
VIII-F3 Privacy
Edge networks are usually closely surrounding people and the data collected by edge devices may include personal and sensitive information of users such as geographical traces, health and activities records, electricity usage, etc. Under the requirement of privacy protection policies, e.g., the European Union’s General Data Protection Regulation (GDPR), sharing and processing of these data across edge service providers may incur privacy leakage and unexpected legal risks. This leads to research questions on privacy-preserved EGI model computation, where cryptology tools like differential privacy and homomorphic encryption can be leveraged.
IX Conclusion

### v2

@@ IX Open Challenges and Future Research Opportunities
IX-A New Promising Applications and Optimizations
@@
IX-A1 New EGI Applications
Besides applications discussed in Sec. V, there are still many burgeoning edge scenarios well matching EGI’s capability. For instance, in digital twin networks, the mapping between physical objects and digital twins can be regarded as graphs and EGI may be applied for task offloading, power allocation, and network slicing [349, 350, 255]. Satellite-aerial-ground intergaetd networks are also one of the emerging edge networks [351, 352, 353], where high-altitude platform stations and ground stations constitute vertices with their communication channels as links. EGI thus can also be used for them.
IX-A2 B5G and 6G Protocols Design
@@
Communication in edge networks, particularly for cross-device services, often involves complex interactions that can be naturally represented as graphs. The networking topology, defined by communication nodes, forms the vertices, while relevant data such as channel capacity, queuing delays, and link states can be modeled as graph properties. This representation opens up opportunities to apply GI for developing advanced communication protocols, such as those for B5G and 6G networks, enabling smarter, faster, and more energy-efficient data transmission. For instance, GNNs can be utilized to dynamically optimize routing strategies and detect intermittent anomalies within edge networks [354, 355]. Furthermore, communication channel states can be incorporated as link attributes in graph models, allowing GI methods to enhance network management and resource allocation [245, 242, 246, 247].
IX-A3 Personalized Services
@@
End users nowadays are immersively surrounded by a vast amount of IoT devices like mobile smartphones, wearable kits, and smart home assets. For individual users, we can weave a user-centric network [356] by collecting all user-related data from these surrounding devices and connecting them in a logic graph, on which EGI techniques can be applied to process, understand, and learn. In other words, EGI can serve as a tool to summarize data points scattered around individual users and be leveraged as a core component to build everyone-centric customized services.
IX-B Edge-Centric Graph Learning Models and Systems
@@
Many existing GI models are originally designed to improve accuracy on specific tasks. In real deployment, however, a set of SLOs beyond accuracy are required by service providers such as responsive latency, system energy, and data privacy, and implementing GI models at edge faces unique challenges. Note that the graphs in edge networks may not be as huge as social networks on Twitter [357]. However, it does not necessarily mean that training and deploying GNN on edge networks is technically trivial.
IX-B1 Resource Constraints to GI Models
@@
As we discussed in Sec. VI, edge devices typically have limited but heterogeneous computational capabilities compared to cloud servers, possessing platforms from MCUs to mobile smartphones to edge servers. Edge-centric GI models need to be lightweight, have acceptable memory footprints, and minimize computational complexity while still delivering accurate results, especially for EGI-based services with real-time or near real-time requirements. This directs innovations in both model level and system level, e.g., lightweight algorithms like computation deduplication [358] and memory-efficient on-device systems [103]. Besides, fine-grained characteristics of GI models also expect precious profiling for refined resource utilization.
IX-B2 System Dynamics
@@
There are two types of system dynamics, namely resource dynamics and graph dynamics. On the one hand, edge devices frequently operate in dynamic and unpredictable resource conditions, including intermittent connectivity, dynamic computing capabilities, or fluctuated communication bandwidths [359, 360]. A stable, satisfactory EGI system is obliged to be robust to these variations and capable of adapting to different network conditions to ensure reliable and timely analysis, especially for those services that process graph data across multiple edge devices. On the other hand, many real-world graph data can be dynamic in both properties and topologies, requiring both algorithmic and system flexibility to adapt to the input variation. Some work [211, 212] exploit online incremental adjustment for minor input variation, where drastic changes like graph reconstitution are still under exploration.
IX-B3 Scaling to Large Graphs
@@
Many real-world networks represented as graphs can be massive and evolve in size, but how to scale EGI systems to large and heavy-attributed graphs poses significant challenges on both the algorithm and system sides. From the algorithm perspective, large graph models borrow wisdom from LLM and incorporate pre-training methods to embrace graphs on giant scales, but also bring overload stress to edge networks. Distributed edge collaboration (cf. Sec. VI-B), which aggregates multiple edge devices as a resource pool for model computation, can be a potential solution but still requires tailored designs with parallel processing, graph partitioning, and incremental learning, etc [361]. From the system perspective, orchestrating a tremendous volume of graph data through edge network channels tackles the communication bottleneck as well as the storage capacity of edge devices. Data compression like graph pruning and feature quantization can be helpful, but how to balance the tradeoff between model accuracy and execution latency requires careful design.
IX-C Emerging Learning Paradigms
@@
IX-C1 Large Graph Model
@@
LGMs borrow the wisdom of LLMs to the graph domain that simultaneously scales both training data and models. This scaling mechanism envisions LGMs to be capable of capturing complex relationships and patterns within large-scale graphs, which are common representations of data in various domains such as social networks, biological networks, and transportation systems [57, 59]. Nonetheless, applying LGM in edge networks severely suffers from the constrained computing capability and network capacity. Edge-cloud collaboration, which marries the local data processing ability of edge networks and the powerful computing power of the cloud, can be a potential solution for deployment. Given the promising ability of large models, LGMs desire more effort in the context of EGI. On the other hand, EGI may also explore the translation between graphs and natural language to exploit the capability of LLMs, e.g., by semantically converting a graph into a paragraph of sentences for LLM input and generation [58].
IX-C2 Learning with Small Data
@@
Edge devices can collect a large volume of data but only with a small portion of labels, which demands EGI to be capable few-shot learners. Towards that, few-shot learning techniques can be promising, which aims at training GL models for accurate predictions on unseen classes or categories with limited labeled data. For example, in some EGI-based object detection tasks [164], GI models can learn to recognize novel objects with a few labeled examples by leveraging information from similar objects and their relationships in a graph representation. Another potential path is transfer learning, which borrows knowledge learned from a source domain or task to improve performance in a target domain or task. An instance of transfer graph learning is anomaly detection in networked systems [264, 265], where GI can learn patterns of fraud by pre-training on a large graph and transfer this knowledge to identify fraudulent activities in a target domain with limited labeled data. TL can also be used for STGNN streams, e.g., for traffic prediction in smart cities [362].
IX-C3 Continual Learning
@@
Continual Learning (CL) for GI is to incrementally retrain GNN models to adapt to new data or tasks without forgetting previous knowledge. It also adjusts the learned latent distribution like Transfer Learning but emphasizes learning patterns from the additional incoming data while preserving previously learned representations. For EGI services, CL enables incremental service refinement for evolved graphs, e.g., in temporal graph data analysis tasks, without training GI models from scratch. Examples include predicting traffic conditions in a road network and running the status of power grids, where continual learning of GI models allows the model to adapt over time and incorporate new temporal information while retaining the knowledge learned from past time steps.
IX-D Native Edge Support for Graph Intelligence
@@
IX-D1 Graph Data Collection
@@
EGI applications can not be realized without pervasive graph data. For many applications, however, these graph data are generated and distributed in different places, e.g., traffic sensory data scattered in road networks and perception information gathered by vehicles in robotic swarms. How to collect these (geographically) distributed data over the edge networks efficiently, i.e., in a high-quality, low-latency, and privacy-preserved manner, thereby becomes an unavoidable problem in running GI-based services. In certain situations, graph generation techniques [363] may be a useful supplement for graph data requirements.
IX-D2 Edge-Cloud Collaboration
@@
Edge networks render computing power in versatile forms, e.g., by a single edge device, a pool of collaborated ones, and even a synergetic edge-cloud continuum. Provisioning these versatile resources efficiently for EGI applications not only requires flexible resource management but also judicious workload scheduling that well aligns with computation and communication. In particular, edge-cloud collaboration provides a combined mechanism to embrace both data locality at the edge and resource richness in the cloud, which allows GI to be more efficient and sustainable and attracts growing attention from the community. Possible means towards this objective include designing GI-oriented communication protocols, GI-specific computation scheduling mechanisms, and a holistic edge network cost optimization methodology.
IX-D3 System Deployment
@@
Given the broad use cases of EGI, their development yet lacks a comprehensive framework with full-stack toolkits, particularly for distributed edge collaboration and edge-cloud synergy. This requires addressing the complexity of distributed systems, including edge device coordination, data synchronization, and fault tolerance, and also tackling long-term interoperability and testing issues. Existing virtualization techniques such as container and function computing can be further explored for friendly EGI development and deployment. The communication aspects, such as timely data transmission of (potentially distributed) graph data and graph modeling of network channels, also desire tailored consideration.
IX-D4 Supporting Large Models
As an edge-oriented methodology, EGI can support large models in the following angles. First, EGI provides an effective way to abstract edge data in the form of graphs, acting as a graph data processing module for large model deployment at the edge. For instance, we may use EGI to transform graph data into prompts acceptable by large language models, allowing language intelligence on edge graphs. Second, with EGI models, edge data can be abstracted into graph embeddings. These embeddings can be used as a knowledge base for multi-modal foundation models aiming at specific edge scenarios, i.e., enabling retrieval-augmented generation. Third, EGI can also be operated as an optimization tool to orchestrate resource allocation for large model computation. To deploy resource-hungry large models on edge platforms, we may utilize many devices in edge networks, where EGI can be a scheduler, as discussed in Sec. VII-A, to schedule efficient resource allocation for large model acceleration.
IX-E Explainability, Security, and Privacy
@@
The open nature of EGI requires trustworthiness across different entities in the ecosystem to make EGI happen. Nonetheless, in many scenarios, edge services are practical only if they are reliable. For EGI, this relates to explainability, security, and privacy.
IX-E1 Explainability
@@
IX-E2 Security
Security holds a dual significance in the context of EGI services. From a model perspective, GI models must be both secure and robust, as they are particularly vulnerable to attacks when malicious vertices are present in the graph input. The risk is further amplified when input graph data is aggregated from multiple entities, increasing the likelihood of fraudulent activity, and underscoring the importance of developing and implementing secure GI models. From a system perspective, edge networks are also susceptible to malicious devices, which can disrupt the learning process or compromise data integrity. This calls for the establishment of verifiable trust mechanisms among edge networks, alongside the development of secure tools to defend against and recover from such attacks.
IX-E3 Privacy
Edge networks typically operate in close proximity to individuals, collecting data from edge devices that may include personal and sensitive information, such as geographic location, health and activity records, and electricity consumption patterns [360]. In light of privacy protection regulations, such as the European Union’s General Data Protection Regulation (GDPR), the sharing and processing of this data across edge service providers poses significant risks of privacy leakage and unforeseen legal liabilities. Consequently, this raises critical research questions regarding not only secure but privacy-preserving computation in EGI models, where cryptographic techniques such as differential privacy and homomorphic encryption can be utilized to mitigate these risks.
X Conclusion